# Integrating Deep-Learning-Based Magnetic Model and Non-Collinear Spin-Constrained Method: Methodology, Implementation and Application

### v2

@@ III.1 Computational Details
All calculations in this work were performed with ABACUS, utilizing plane waves or numerical atomic orbitals Li et al. (2016) to expand wavefunctions. The pseudopotentials were sourced from the Pseudo-Dojo pseudopotential library Pseudo Dojo . The numerical atomic orbitals of ABACUS are generated by optimizing the spillage function averaged over structures; this strategy was originally proposed in the works of Chen et al. Chen et al. (2010) (denoted as v1.0) and improved by Lin et al. Lin et al. (2020) by including the kinetic terms (denoted as v2.0). In this work, a numerically further improved version is also used (denoted as v2.1). These v2.1 orbitals have not yet been officially released on the official website but the generation code is online available orb . This paper performs NAOs calculations using v2.0 and v2.1 orbitals, including Double-$\zeta$ plus polarization functions (DZP), and Triple-$\zeta$ plus double polarization functions (TZDP) with different cutoffs. Furthermore, for all calculations, the cutoff energy was set to 100 Ry, and the Brillouin zone was uniformly sampled at 0.14 Bohr-1 intervals. The projected orbitals used to calculate magnetic moments were modified numerical atomic orbitals, with a modulation radius of 3 Bohr. Each inner loop of $\bm{\lambda}$ optimization converges when the maximum atomic magnetism variation is less than $\delta\mathbf{M}<10^{-7}\mu_{B}$, while the SCF convergence until density difference $\delta\rho<10^{-6}$.
III.2 Projection Methods
As mentioned in Sec. II.2, we have developed an innovative algorithm designed to estimate and control atomic magnetic moments using localized orbital projection techniques. The key innovation of our method lies in the localized modulation of the numerical atomic orbital basis. We have benchmarked this algorithm on ferromagnetic (FM) and antiferromagnetic (AFM) iron (Fe) bulk.
Figure 3: Estimation of BCC-Fe atomic magnetic moments as a function of modulation radius $r_{m}$ using the modulated orbital projection algorithm, employing DZP and TZDP basis sets with cutoff radius $r_{c}$ ranging from 6.0 to 10.0 Bohr, to demonstrate the impact of different NAO basis sets on atomic magnetic moments. Where (a) is for the PW basis set in the FM magnetic configuration, with the reference value being the total magnetization per atom (TMAG); (b) is for the NAO basis set in the FM magnetic configuration, with the reference value being the TMAG; (c) is for the PW basis set in the AFM magnetic configuration, with the reference value being the absolute magnetization per atom (AMAG); and (d) is for the LCAO basis set in the AFM magnetic configuration, with the reference value being the atomic magnetization from Mulliken charge.
As shown in Fig. 3(a) and (c), the projected magnetic moments obtained from the different modulated orbitals agree well with each other if $r_{m}<4$ Bohr in all PW calculations. This indicates that the projection results are dependent solely on the selected $r_{m}$, rather than on $r_{c}$. The results from the NAO basis (Fig. 3(b) and (d)) show differences in the projection results between various basis sets, with these discrepancies arising solely from the NAO basis set itself. Furthermore, the comparison between left and right panel of Figs. 3 shows the higher-precision TZDP results align well with the PW ones, while the DZP results ensure qualitative correctness.
The choice of an appropriate modulation radius $r_{m}$ significantly affects the estimated values of atomic magnetic moments. We discuss two possible strategies for determining $r_{m}$. The first approach is to calibrate the projected orbital radius by matching it to reference values, such as experimental data or literature values. However, this can be challenging for complex compounds or systems without available experimental data, as it may be difficult to identify suitable reference values. The second approach is to select the modulation radius corresponding to the maximum atomic magnetic moment. This is based on the rationale that, at this point, the projected orbitals exhibit the largest difference in projection results of wave functions with different spins, thereby providing the most accurate description of atomic magnetic moments and the energy variation caused by spin.
For BCC-Fe, the modulation radius chosen using the first approach is 2.0 Bohr, with the FM and AFM atomic magnetic moments being 2.21 and 1.48 $\mu_{B}$, respectively. These values are in good agreement with those calculated by VASP Ma and Dudarev (2015), which are 2.22 and 1.52 $\mu_{B}$. The second approach determines the radius to be 3.0 Bohr, with the estimated atomic magnetic moments being 2.36 $\mu_{B}$ for FM and 1.6 $\mu_{B}$ for AFM. In this work, we adopted the more general second strategy, selecting 3.0 Bohr as the target radius. Subsequent tests will demonstrate that this choice of projected orbital provides a reliable description of the magnetic potential energy surface and related physical quantities.
III.3 Finite Difference Tests
Figure 4: Finite difference tests for (a) the atomic force in BCC Fe16, (b) the magnetic torque in ternary alloy NiMnTi (c) the cell stress in binary alloy FeNi. Here “Analytic” represents the value calculated from the formula, while “Numerical” is the value determined from the finite difference method. “11” refers to the component of stress matrix $\sigma_{11}$.
Based on the theory defined in Section II, we implemented the spin-cDFT method in the open-source software ABACUS, the method is available with either plane wave or numerical atomic orbital basis sets. In the spin-cDFT framework, we introduced corrections for the energy, atomic forces, and lattice stresses under full magnetic constraints and incorporated calculations of magnetic torque $\bm{\lambda}$. These magnetic torques are optimized in the inner loop of the self-consistent field iterations. To validate our implementation, we compared analytical solutions for atomic forces, lattice stresses, and magnetic torques against numerical ones obtained through finite difference method. As shown in Fig. 4, these tests were conducted on elemental body-centered cubic (BCC) iron (Fe), FeNi binary alloy, and NiMnTi ternary alloy. For a 16-atom BCC-Fe supercell, we perturbed one Fe atom’s position, resulting in a close match between the finite-difference derived forces and the analytical values. In the case of NiMnTi, we perturbed an atom and found high consistency between analytical forces and finite-difference stress tensor values. Crucially, magnetic torques are the partial derivatives of the energy with respect to magnetic moments, rather than atomic positions. Selecting one Fe atom in the FeNi alloy, we perturbed its magnetic moment, which showed perfect agreement between energy-derived magnetic forces and analytically obtained values from the inner loop. For these three tests, the cutoff energy was set to 100 Ry. The Brillouin zone was uniformly sampled at 0.14 Bohr-1 for Fe16 and 0.15 Bohr-1 for FePt/NiMnTi. The successful finite difference tests corroborate the proper functionality of the cDFT features for energy, atomic forces, magnetic torques, and lattice stress in ABACUS.
III.4 BCC and FCC Structures of Iron
Figure 5: Comparison between the unconstrained and constrained DFT results for ferromagnetic BCC-Fe (a) and FCC-Fe (b). The solid lines show the total energy per atom as a function of the cell volume per atom. Here “DZP-8au-v2.1” represents the v2.1 NAOs calculations based on the DZP orbitals with the 8 Bohr cutoff. The gray histograms represent the atomic magnetic moment $m$ after fully unconstrained self-consistent calculations. The dotted lines show the cDFT energy with certain constrains, and plot the corresponding magnetic force $\lambda=|\bm{\lambda}|$ in the upper panel.
The BCC and FCC phases are two prevalent iron crystal structures. The BCC structure is the stable phase of iron at room temperature and is known for its high strength and low ductility. The BCC Fe structure exhibits a ferromagnetic (FM) state with atomic magnetism around 2.2 $\mu_{B}$ per atom Acet et al. (1994), and the Curie temperature was found at 1043 K experimentally Chen and Sundman (2001). On the other hand, The FCC structure has a higher density of atoms compared to BCC. Experimentally, iron undergoes a structural phase transition from BCC to FCC around 1085 K Okamoto et al. (2016) and the FCC structure is the stable phase until 1667 K Okamoto et al. (2016). Although FCC Fe exhibits paramagnetic behavior in the experiment, theoretically, it is predicted that the energy of either antiferromagnetic (AFM) or double-layer antiferromagnetic (DAFM) states would be lower than that of ferromagnetic at 0 K Rinaldi et al. (2024). It has a more unimodal density of states at the Fermi level, which results in lower magnetic moments compared to BCC Fe.
Figure 6: Total energy of BCC-Fe (a) and FCC-Fe (b) as a function of atomic magnetic moment and cell volume. All calculations are performed by spin-cDFT with DZP-8au-v2.1 NAOs orbitals.
Figure 7: (a) the BCC-Fe with 2 Fe atoms labels as $\{M_{0}-\Delta M,\theta\}$, where the two atoms have the same atomic magnetic moment $M_{0}-\Delta M$. $M_{0}$ is the atomic magnetic moment and $\Delta M$ is ranging from -0.4 $\mu_{B}$ to 0.4 $\mu_{B}$. $\theta$ is the angle between them. (b-c) shows the magnetic energy barrier difference between the NAOs basis and PW basis, namely $\Delta\hat{E}(M,\theta)=\hat{E}(M,\theta)|_{\rm NAOs}-\hat{E}(M,\theta)|_{\rm PW}$. Here the magnetic energy barrier is defined by the energy above the ground state $\hat{E}(M,\theta)=E(M,\theta)-E(M_{0},0)$. (d-f) shows the ratio of the magnetic energy barrier difference $\Delta\hat{E}(M,\theta)$ to the PW results.
To systematically present the potential energy surfaces of Fe, we modeled BCC-Fe and FCC-Fe with atomic volumes ranging from 8 to 16 Å3 and calculated their total energies at various magnetic moments using spin-cDFT. The magnetic moments considered ranged from 0.2 to 3.8 $\mu_{B}$. Fig. 6 illustrates the potential energy surfaces, $E(V,M)$, for BCC-Fe and FCC-Fe, respectively. The results reveal that the BCC phase exhibits a single energy minimum, whereas the FCC phase features two distinct energy minima, consistent with the findings presented in Fig. 5.
Magnetic calculations have high accuracy requirements. Unlike the PW basis, where accuracy can be systematically improved by increasing the number of plane waves, the accuracy of the NAOs basis depends on the quality of the orbitals themselves and lacks a systematic way to enhance precision. In the following, we will quantitatively examine the precision of spin-cDFT results using different basis sets. In the Heisenberg model $H=\sum_{ij}J_{ij}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, the magnetic exchange strength $J$ determines the resistance of the material’s magnetic order to external perturbations such as temperature and magnetic fields. This strength typically falls within the range of a few to tens of meV Szilva et al. (2023). The value of $J$ can generally be derived from the energy differences between various magnetic configurations, making the magnetic barrier energy an excellent quantitative metric for evaluating accuracy.
Fig. 7 presents the results for various NAOs basis sets, showing the energy error compared to the PW results as $\Delta\hat{E}(M,\theta)=\hat{E}(M,\theta)|_{\rm NAOs}-\hat{E}(M,\theta)|_{\rm PW}$. According to the figure, except in regions where the magnetic moment decreases and the angle is relatively large—where the energy error is significantly higher—the error is fairly uniform in other areas. The results indicate that the error primarily depends on the type of basis rather than the cutoff radius. In practical scenarios, such as the gradual transition from a ferromagnetic to a nonmagnetic state with increasing temperature in BCC-Fe, the magnetic moments significantly change in angle instead of amplitude. When the angle changes from $0^{\circ}$ to $180^{\circ}$, the error for the DZP basis set ranges from 2 to 5 meV, whereas the error for the TZDP basis set is significantly reduced to 0.5 to 1 meV. The experimental Curie temperature of BCC-Fe is 1043 K, corresponding to a thermodynamic energy of approximately 90 meV. Given this, both DZP and TZDP basis sets are expected to provide relatively accurate $J$ values, with the precision of the TZDP basis set nearly matching that of the PW basis.
To validate this inference within the framework of density functional theory, we analytically calculated the magnetic exchange energy using the magnetic force theorem Liechtenstein et al. (1984, 1987). Specifically, we utilized the exchange energy calculations by ABACUS and TB2J He et al. (2021), employing a $9\times 9\times 9$ supercell to determine the $J_{ij}$ values up to the fourth-nearest neighbors. The results obtained using different NAOs basis are summarized in Table 1. Since ABACUS currently does not support an interface with PW and TB2J, we used PW results reported in the literature as reference values Wang et al. (2010); Pajda et al. (2001). It is evident that the v2.1 NAOs orbitals are more accurate than those in version 2.0, and the results obtained with the TZDP basis set closely match the reference values, with an error margin within 5%.
Table 1: Exchange energy of ferromagnetic BCC-Fe calculated by various NAOs basis. The cutoff radii is 7 Bohr for all orbitals. “NN” represents the nearest-neighbor exchange, “2NN” refers to the second one, and beyond.
$J_{ij}$ (meV) DZP-v2.0 DZP-v2.1 TZDP-v2.1 Ref. Wang et al.,2010 Ref. Pajda et al.,2001
NN 11.73 17.88 20.23 16.57 19.48
2NN 12.06 11.58 12.23 14.69 11.09
3NN -0.05 -0.36 -1.11 -0.57 -0.20
4NN -2.17 -1.15 -1.93 -2.52 -1.71
III.5 DeePSPIN Model
Figure 8: For (a) BCC and (b) FCC Fe, the comparison of the energy predicted by the DeePSPIN model with those calculated by ABACUS cDFT for the collinear magnetic configurations (FM/AFM/DAFM) with various atomic magnetic moments.
The non-collinear spin-cDFT implementation in ABACUS provides a powerful tool for studying complex magnetic configurations at the atomic scale. However, conventional DFT methods are time consuming, which makes it challenging to simulate magnetic dynamic processes, such as transitions in magnetic ordering that require large-scale supercells. DeePSPIN Yang et al. (2024) is a deep learning approach for magnetic materials that treats spin as so-called “pseudo-atoms” and integrates with the descriptor framework of DeepPot-SE Zhang et al. (2018), preserving translational, rotational, and permutation symmetries. In detail, the DeePSPIN model requires high-precision first-principles data for magnetic materials, including energy, forces, magnetic torques, stress (optional), and atomic configurations, as training data. Additionally, the design of appropriate loss functions and active learning methods Zhang et al. (2020) significantly reduces the required number of samples. A well-trained DeePSPIN model can accurately predict physical properties such as energy, forces, and magnetic torques. By combining it with methods like molecular dynamics (MD) Thompson et al. (2022) and the Landau-Lifshitz-Gilbert equations Gilbert (2004), it can simulate spin evolution in large-scale systems and accurately describe spin-lattice interactions.
Figure 9: For ferromagnetic (a) BCC and (b) FCC Fe, the energy predicted by the DeePSPIN model as a function of volume with various atomic magnetic moments.
To assess the capability of ABACUS+DeePSPIN in representing significant changes in the lattice and magnetic moments, we generated over 10,000 collinear magnetic configurations for the BCC/FCC phase of iron, including three magnetic states of ferromagnetic (FM), antiferromagnetic (AFM), and double-layer ferrimagnetic (DAFM). The cell volume perturbations ranged from -30% to +30%, and the magnitude of the atomic magnetic moments varied from 0 to 4.0 $\mu_{B}$. These data points were obtained through collinear spin-cDFT calculations using the DZP-8au-v2.1 basis set. We then trained a DeePSPIN model based on these perturbed configurations. Fig. 8 adopts both DeePSPIN and DFT methods to predict the total energy of collinear magnetic configurations such as FM, AFM, and DAFM. We find the results from the two models align well with each other, demonstrating the good accuracy of the DeePSPIN model. Fig. 9 illustrates the dependence of total energy for varying magnitudes of magnetic moments, with the volume being changed isotropically. The solid line represents the ground-state energy from unconstrained DFT calculations at each volume, while the data points from the DeePSPIN model correspond to excited states, which have energies higher than those of the DFT results. These observations are in good agreement with recent studies Rinaldi et al. (2024), demonstrating the reliability of the ABACUS spin-cDFT data and the expressive capability of the DeePSPIN model.
Figure 10: Magnetization of BCC-Fe as a function of temperature. Based on the DeePSPIN model, MD simulations at different temperatures are performed, with the stable magnetic moment counted as the magnetization at that temperature. Line width refers to the error bar. The number of atoms in the supercell is 1024, and the virtual magnetic mass M=0.01.
To explore more complex magnetic phenomena, such as the transition of magnetic ordering, it is crucial for the model to accurately capture information about non-collinear magnetic moments. To achieve this, we used ABACUS to generate first-principles data and trained a DeePSPIN model for iron using an active learning strategy Zhang et al. (2020). In detail, we performed random perturbations of the unitcell, atomic positions, and magnetic moments based on the ground-state configurations of BCC-FM/FCC-AFM/DAFM ($2\times 2\times 2$ supercell including 16 Fe atoms). For each configuration, we applied 100 random perturbations, which were used as the initial training data for DPGEN sampling. The maximum perturbation amplitude for the unitcell was 1%, while the maximum perturbation for atomic positions was 0.03 Å. Magnetic moments were randomly rotated by up to 90 degrees, and their magnitudes were perturbed by up to 0.5 $\mu_{B}$. Sampling was conducted separately for BCC and FCC configurations. The sampling temperature ranged from 50 to 1600 K, with each temperature range including MD configurations with the virtual magnetic mass of 0.01/0.05/0.1/0.5/1.0. We employed both NAOs basis (DZP-7au-v2.0) and PW basis to generate two separate data sets. All computations were performed using non-collinear spin-cDFT, incorporating spin-orbit coupling. The PW training strategy followed the same approach as for the NAOs, with the key difference being that NAOs BCC sampling involved 400 configurations per round, whereas PW BCC sampling involved 200 configurations per round. Ultimately, the NAOs basis set generated a total of 34,703 DPGEN samples, while the PW basis set generated 20,931 DPGEN samples. Additionally, for each basis set, we supplemented the data with 175 perturbed configurations for FCC-FM and 109 configurations for the FCC $4\times 2\times 2$ supercells. It is worth noting that DPGEN employs a custom-built version of LAMMPS, designed to be compatible with DeepSpin et al. (2025), for NVT sampling.
We trained two models, DeePSPIN-DZP and DeePSPIN-PW, based on first-principles data, including energy, force, and magnetic torque. These models allow us to perform large-scale simulations to observe the evolution of magnetism with temperature. We conducted NVT ensemble simulations at different temperatures using the trained models, employing an $8\times 8\times 8$ supercell (consisting of 1024 atoms), with a virtual magnetic mass set to 0.01, a timestep of 0.1 fs, for a total of 3 ps of simulation. After the MD equilibration, we selected the trajectory from 1 to 3 ps to calculate the average total magnetization of the system. In Fig.10, we present the total magnetization of BCC-Fe as a function of temperature. It can be observed that BCC-Fe exhibits ferromagnetic behavior at low temperatures, with the total magnetization exceeding 2 $\mu_{B}$. The magnetization decreases gradually up to 800 K, before undergoing a sudden drop around 1000 K, at which point it approaches zero, indicating a ferromagnetic-to-paramagnetic transition. This transition temperature is in close agreement with the experimentally observed Curie temperature of 1043 K. Surprisingly, although the DZP-7au-v2.0 basis set introduces non-negligible errors when calculating the magnetic exchange strength $J$, the magnetization as a function of temperature and the Curie temperature is in very close agreement with the results from PW basis.
IV Conclusion
The DeePSPIN model trained with the first-principles data generated by this method can predict the energy for different volumes and magnetic moments with high accuracy, demonstrating its good adaptability. Combined with molecular dynamics simulations, the DeePSPIN model successfully observed the ferromagnetic-paramagnetic transition near the experimental Curie temperature.
In conclusion, the noncollinear spin-constrained method implemented in ABACUS provides a powerful tool for studying complex magnetic phenomena at the atomic scale, and a data engine for deep-learning magnetic models. Its integration with AI-assisted magnetic models enables large-scale magnetic simulations, providing new possibilities for the study of complex magnetic phenomena.
Acknowledgements.

### v3

@@ 4.4 First-principles Calculations
All calculations in this work were performed with ABACUS, utilizing plane waves or numerical atomic orbitals [Li2016] to expand wavefunctions. The pseudopotentials were sourced from the Pseudo-Dojo pseudopotential library [PseudoDojo]. The numerical atomic orbitals of ABACUS are generated by optimizing the spillage function averaged over structures; this strategy was originally proposed in the works of Chen et al. [Chen2010] (denoted as v1.0) and improved by Lin et al. [lin_accuracy_2020] by including the kinetic terms (denoted as v2.0). In this work, a numerically further improved version is also used (denoted as v2.1).
These v2.1 orbitals have not yet been officially released on the official website but the generation code is online available [orb-v2.1]. This paper performs NAOs calculations using v2.0 and v2.1 orbitals, including Single-$\zeta$ (SZ), Double-$\zeta$ plus polarization functions (DZP), and Triple-$\zeta$ plus double polarization functions (TZDP) with different cutoffs. The convergence test for the plane-wave cutoff energy is presented in the Supplementary Information. By adopting a criterion of energy error per atom below 1 meV and stress error below 0.1 kbar, the cutoff energy was set to 100 Ry for all calculations (see Fig. S4). The Brillouin zone was uniformly sampled at 0.14 Bohr-1 intervals. Specifically for the BCC-Fe supercell with 16 iron atoms, this k-spacing corresponds to a $5\times 5\times 5$ Monkhorst-Pack grid. The projected orbitals used to calculate magnetic moments were modified numerical atomic orbitals, with a modulation radius of 3 Bohr. Each inner loop of $\bm{\lambda}$ optimization converges when the maximum atomic magnetism variation is less than $\delta\mathbf{M}<10^{-7}\mu_{B}$, while the SCF convergence until density difference $\delta\rho<10^{-6}$. All noncollinear calculations incorporated the spin-orbit coupling (SOC) effect. We would like to clarify that while SOC effects in pure Fe are relatively weak—smaller than the intrinsic errors of our model—we deliberately retained SOC calculations to rigorously test the stability of our implementation under combined noncollinear magnetism, SOC, and cDFT conditions, as well as to validate the robustness of the entire model workflow. The cDFT method implemented in ABACUS, capable of precisely obtaining first-principles results for arbitrary magnetic moment configurations, also serves as an effective tool for studying the systems where SOC plays a decisive role in determining the magnetic ground state or emergent phenomena [2023-yang-first]—such as heavy-element magnets, topological spin textures, or Dzyaloshinskii–Moriya interaction (DMI) driven systems. In Table S4 of the Supplementary Material, we have included a short example using the four-states method [2011-xiang-predicting] to calculate the DMI effect in the Fe–Pt system, demonstrating this capability.
4.5 Molecular Dynamics Simulation
The molecular dynamics simulations utilize a modified version of LAMMPS using the TSPIN method involved simultaneous integration of lattice and spin degrees of freedom within a unified Nosé–Hoover chain (NHC) thermostat framework [2025-huang-prep]. The dynamics are propagated using symplectic integration schemes derived from an extended Hamiltonian formulation, which preserves symplecticity and ensures accurate sampling of canonical (NVT) and isothermal–isobaric (NPT) ensembles.
The Lagrangian that governs the system’s evolution can be expressed as:
$L=\sum_{i}\frac{p_{si}^{2}}{2\mu_{i}}+\sum_{i}\frac{p_{i}^{2}}{2m_{i}}-U(\{\mathbf{R}\},\{\mathbf{M}\}),$ (43)
where $p_{si}$ and $p_{i}$ are the generalized momenta associated with the spins and the lattice positions, respectively. $\mu_{i}$ denotes the virtual magnetic mass. The energy function $U(\{\mathbf{R}\},\{\mathbf{M}\})$ accounts for the interactions between the lattice, spins, and spin-lattice couplings. The Euler-Lagrange equations derived from this Lagrangian provide the following equations of motion:
$\dot{\mathbf{R}}_{i}$ $=\frac{p_{i}}{m_{i}},\quad$ $\dot{p}_{i}=-\frac{\partial U}{\partial\mathbf{R}_{i}}=\mathbf{F}_{i},$ (44)
$\dot{\mathbf{M}}_{i}$ $=\frac{p_{si}}{\mu_{i}},\quad$ $\dot{p}_{si}=-\frac{\partial U}{\partial\mathbf{M}_{i}}=\bm{\lambda}_{i}.$
The above equations describe the evolution of the lattice ($\mathbf{R}_{i}$) and spin ($\mathbf{M}_{i}$), where the lattice evolves according to the atomic forces ($\mathbf{F}_{i}$), while the spins evolve under the magnetic torque ($\bm{\lambda}_{i}$). For the NVT and NPT ensembles, we use the equations of motion with thermostat variables for temperature control shown in Ref. [2025-huang-prep].
4.6 Finite Difference Tests
For the force test, we randomly selected one iron atom from a BCC-Fe16 configuration and perturbed its z-coordinate with a step size of $\Delta=0.01~\text{Bohr}$, calculating the energies of 11 perturbed configurations $E(z_{0}+i\Delta)$. Using the finite difference formula:
$\mathbf{F}_{z}(i)=\frac{\partial E}{\partial\mathbf{r}_{z}}\approx\frac{f((i+1)\Delta)-f(-(i-1)\Delta)}{2\Delta},$ (45)
we computed numerical solutions for 9 points (excluding the edge points).
For the magnetic torque test, we randomly selected a FePt configuration and perturbed the z-component magnetic moment of one Fe atom with a step size of $\Delta=0.1~\mu_{B}$. We calculated the total energies of 11 perturbed configurations and determined the numerical solutions $\frac{\partial E}{\partial\mathbf{m}_{z}}$ using the central difference method.
We randomly selected a NiMnTi configuration to conduct stress finite-difference testing. The stress can be obtained through strain perturbation:
$\sigma_{\alpha\beta}=-\frac{1}{\Omega}\frac{\partial E}{\partial\varepsilon_{\alpha\beta}},$ (46)
where $\Omega$ represents the volume, and the stress tensor $\sigma$ has six independent components: $\sigma_{11}$, $\sigma_{12}$, $\sigma_{13}$, $\sigma_{22}$, $\sigma_{23}$, $\sigma_{33}$. For each component, we constructed corresponding strained configurations. Taking $\sigma_{11}$ as an example, we generated 11 configurations with $\varepsilon_{11}=i*\Delta$ (step size $\Delta=0.0001$). According to the following formula, the strain induces the corresponding unit cell modifications:
$\varepsilon_{11}:\begin{pmatrix}a_{11}+\varepsilon_{11}a_{11}&a_{12}&a_{13}\\ a_{21}+\varepsilon_{11}a_{21}&a_{22}&a_{23}\\ a_{31}+\varepsilon_{11}a_{31}&a_{32}&a_{33}\end{pmatrix}$ $\varepsilon_{12}:\begin{pmatrix}a_{11}+\varepsilon_{12}a_{12}&a_{12}&a_{13}\\ a_{21}+\varepsilon_{12}a_{22}&a_{22}&a_{23}\\ a_{31}+\varepsilon_{12}a_{32}&a_{32}&a_{33}\end{pmatrix}$ (47)
$\varepsilon_{22}:\begin{pmatrix}a_{11}&a_{12}+\varepsilon_{22}a_{12}&a_{13}\\ a_{21}&a_{22}+\varepsilon_{22}a_{22}&a_{23}\\ a_{31}&a_{32}+\varepsilon_{22}a_{32}&a_{33}\end{pmatrix}$ $\varepsilon_{13}:\begin{pmatrix}a_{11}+\varepsilon_{13}a_{13}&a_{12}&a_{13}\\ a_{21}+\varepsilon_{13}a_{23}&a_{22}&a_{23}\\ a_{31}+\varepsilon_{13}a_{33}&a_{32}&a_{33}\end{pmatrix}$
$\varepsilon_{33}:\begin{pmatrix}a_{11}&a_{12}&a_{13}+\varepsilon_{33}a_{13}\\ a_{21}&a_{22}&a_{23}+\varepsilon_{33}a_{23}\\ a_{31}&a_{32}&a_{33}+\varepsilon_{33}a_{33}\end{pmatrix}$ $\varepsilon_{23}:\begin{pmatrix}a_{11}&a_{12}+\varepsilon_{23}a_{13}&a_{13}\\ a_{21}&a_{22}+\varepsilon_{23}a_{23}&a_{23}\\ a_{31}&a_{32}+\varepsilon_{23}a_{33}&a_{33}\end{pmatrix}$
The raw data is shown in the Table. S1/S2/S3 of Supplementary Information.
Data availability
The data and models used in this study have been made open-source and are hosted on AIS Square (https://www.aissquare.com/), where they can be accessed online (https://www.aissquare.com/datasets/detail?pageType=datasets&name=Fe-DeepSpin&id=386).
Code availability
The codes used to produce the results are available from the corresponding author upon reasonable request. The relevant implementation is scheduled to be open-sourced in the next major release of ABACUS and may be accessed at: https://github.com/deepmodeling/abacus-develop.
Author contributions
D.Z. and W.Z. designed and implemented the spin-constrained method in ABACUS. Z.C. participated in the discussion, design, and benchmark of the implementation. Y.H. and L.Z. provided guidance on the projection method. X.P. and Y.W. constructed the first-principles database and trained the iron model. D.Z. and Z.H. implemented code for magnetic model training and molecular dynamics simulations. W.Z. and D.Z. wrote the main manuscript. W.Z., B.X. and M.C. conceived and directed the research, supervised the analysis and interpretation. All authors contributed to the discussions and the final editing.
Competing interests
The authors declare no competing financial or non-financial interests.
Acknowledgements